# Anti-de Sitter $\bm{\to}$ de Sitter transition driven by Casimir forces and mitigating tensions in cosmological parameters

### v2

@@ Abstract
Very recently, an empirical conjecture to simultaneously resolve the $H_{0}$ and $S_{8}$ tensions has been contemplated Akarsu:2023mfb . The conjecture, which is rooted on the graduated dark energy model Akarsu:2019hmw ; Akarsu:2021fol ; Akarsu:2022typ , postulates that $\Lambda$ may have switched sign (from negative to positive) at critical redshift $z_{c}\sim 2$. The so-called “$\Lambda_{s}$ model” then suggests that the Universe may have recently experienced a phase of rapid transition from anti-de Sitter (AdS) vacuum to de Sitter (dS) vacuum. However, the AdS swampland conjecture Lust:2019zwm indicates that AdS vacua are an infinite distance apart from dS vacua in moduli space. Thus, the postulated switch in sign of $\Lambda$ at zero temperature seems unlikely. In this Letter we investigate whether quantum effects derived from the Casimir energy of fields inhabiting the bulk could produce the hypothesized AdS $\to$ dS crossover transition. [^1] For related work, see Ye:2020btb .
Before proceeding, we pause to discuss the benefits of the $\Lambda_{s}$ model. The model simultaneously resolves the major cosmological tensions including: $H_{0}$, $S_{8}$, and SNe Ia absolute magnitude Akarsu:2023mfb . In addition, it can accommodate the BAO Lyman-$\alpha$ disicrepancy Akarsu:2019hmw . [^2] The recent investigation presented in Akarsu:2023mfb is based on: the Planck CMB data Planck:2019nip , the Pantheon+ SNe Ia sample Brout:2022vxf , the data release of KiDS-1000 KiDS:2020suj , and the (angular) transversal 2D BAO data on the shell Nunes:2020hzy ; deCarvalho:2021azj , which are less model dependent than the 3D BAO data used in previous studies of $\Lambda_{s}$. It is important to stress that the BAO 3D data sample assumes $\Lambda$CDM to determine the distance to the spherical shell, and hence could potentially introduce a bias when analyzing beyond $\Lambda$CDM models Bernui:2023byc . Furthermore, for flat cosmology, $\Omega_{m}+\Omega_{\Lambda}\simeq 1$, from which it follows that the mean energy density in the universe is equal to the critical density, where $\Omega_{m}$ is the fractional nonrelativistic matter density and $\Omega_{\Lambda}$ is the cosmological-constant energy density. This implies that $\Omega_{m}>1$ when $\Lambda<0$, in agreement with the otherwise puzzling JWST observations Adil:2023ara . [^3] See however Sabti:2023xwo and Padmanabhan:2023esp .
@@
This potential possesses a minimum at $R=R_{1}$, for which one can show that it always satisfies $R_{1}>R_{0}$. So the radius gets increased due to the uplift, and hence the associated species scale Dvali:2007hz ; Dvali:2007wp ; vandeHeisteeg:2022btw ; Cribiori:2022nke ; vandeHeisteeg:2023dlw decreases (see also the discussion below). It follows that the corresponding species entropy Cribiori:2023ffn always increases due to the uplift.
@@
Note that for $a-a^{c}\to 0^{+}$, the minimum $R_{1}$ is near $R_{0}$, consistently with the quadratic approximation given in (1). Thus, the parametric dependence between the vacuum energy and the radius at its minimum is the same as before the uplift, in accordance with the dark dimension scenario Montero:2022prj . A point worth noting at this juncture is that the relation $|V|=R^{-1/4}$ is at the heart of vacuum energy calculations in string theory, see Anchordoqui:2023laz for a recent discussion.
@@
where $\Lambda_{5}$ is the 5D cosmological constant, $r\equiv\langle R\rangle$, $T_{4}$ is the brane tension, and
$V_{C}(R)=2\pi R\left({r\over R}\right)^{2}\sum_{i}(-1)^{s_{i}}\ N_{i}\ \rho(R,m)$ (9)
corresponds to the Casimir energy, and where the sum goes over all 5D states in the spectrum, $N_{i}$ is the number of degrees of freedom of the $i$-th particle, $s_{i}=0\,(1)$ for bosons (fermions), and the factor $(r/R)^{2}$ comes from the rescaling of the 4D metric in the transformation to the Einstein frame Anchordoqui:2023etp . The Casimir energy density for a particle of mass $m$ (with periodic boundary conditions) is given by
@@
where $K_{5/2}(x)$ is the Bessel function Arkani-Hamed:2007ryu .
@@
$V_{C}(R)=\sum_{i}\frac{2\pi r^{2}}{32\pi^{7}R^{6}}\ (N_{F}-N_{B})\ \Theta(R_{i}-R)\,,$ (12)
where $m_{i}=R_{i}^{-1}$, $\Theta$ is a step function, and $N_{F}-N_{B}$ stands for the difference between the number of light fermionic and bosonic degrees of freedom Anchordoqui:2023etp . Eventually, the fermionic degrees of freedom overwhelm the bosonic contribution, giving rise to possible minima, as long as $R_{i}<R_{\rm max}$.
Figure 1: The potential $V(R)$ for $\Lambda_{5}=(23~{}{\rm meV})^{5}$ and $|T_{4}|=(25~{}{\rm meV})^{4}$ considering $N_{F}-N_{B}=3$ (left) and $N_{F}-N_{B}=7$ (right).
[…]
@@
We can also envision a more complex dark sector in which the gauge field is replaced by a complex scalar with Yukawa coupling to the neutrinos (that could provide them with mass). The spectrum has addition of an extra real scalar, such that before decay of the complex scalar $N_{F}-N_{B}=4$ and after decay $N_{F}-N_{B}=6$. In Fig. 2 we show the corresponding AdS $\to$ dS transition for this scenario. The resulting $\Delta N_{\rm eff}=0.30$ is also consistent with Planck data Planck:2018vyg .
Figure 2: The potential $V(R)$ for $\Lambda_{5}=(23~{}{\rm meV})^{5}$ and $|T_{4}|=(25~{}{\rm meV})^{4}$ considering $N_{F}-N_{B}=4$ (left) and $N_{F}-N_{B}=6$ (right).
The next-to-minimal dark sectors that can accommodate the AdS $\to$ dS crossover transition can be nicely combined with the dark dimension proposal to elucidate the radiative stability of the cosmological hierarchy $\Lambda/M_{p}^{4}\sim 10^{-120}$ Montero:2022prj . As can be seen in Figs. 1 and 2, the connection is possible for a compactification radius $R_{\perp}^{\rm(AdS)}\sim 8~{}\mu{\rm m}$, which implies the Kaluza-Klein (KK) tower of the dark dimension opens up at $m_{\rm KK}^{\rm(AdS)}\sim 155~{}{\rm meV}$. After the vacuum undergoes the AdS $\to$ dS transition the compact space slightly enlarges to $R_{\perp}^{\rm(dS)}\sim 10~{}\mu{\rm m}$ and so $m_{\rm KK}^{\rm(dS)}\sim 123~{}{\rm meV}$. This in turn makes a shift of the species scale,
@@
Another possible course of action to accommodate the AdS $\to$ dS transition (with $N_{F}-N_{B}=4$ in the AdS phase and $N_{F}-N_{B}=6$ in the dS phase) is to assume that the complex scalar $\varphi$ has a potential with two local minima with very small difference in vacuum energy and bigger curvature (mass) of the lower one. More concretely, the potential is given by
@@
where $V_{\varphi}=\partial V/\partial\varphi$. The fourth degree of freedom is adjusted by requiring the height of the barrier to accommodate a decay rate of the false vacuum
@@
where $Z(T)=\int[dx]\ e^{-S_{E}/\hslash}$ is the path integral, $[dx]$ is a functional measure, and $S_{E}$ the Euclidean action Coleman:1980aw ; Hawking:1981fz . If we now take $\Lambda_{5}\sim(23~{}{\rm meV})^{5}$ and $|T_{4}|=(25~{}{\rm meV})^{4}$, the vacuum energy densities of the AdS and dS phases are those shown in Fig. 2.
It is of interest to further investigate the impact of adding extra-relativistic degrees of freedom into the $\Lambda_{s}{\rm CDM}$ model. In particular, a value of $\Delta N_{\rm eff}\neq 0$ would accelerate the universe before the CMB epoch, and hence could modify the critical redshift $z_{c}$ for the AdS $\to$ dS crossover transition. A study along this line is underway using the Boltzmann solver CLASS Blas:2011rf ; Lesgourgues:2011rh in combination with MontePython Audren:2012wb ; Brinckmann:2018cvx (and the latest results from the Planck mission Planck:2019nip , Pantheon+ SNe Ia Brout:2022vxf , BAO Nunes:2020hzy ; deCarvalho:2021azj , and KiDS-1000 KiDS:2020suj ) to perform a Monte Carlo Markov Chain study of the $\Lambda_{s}{\rm CDM}+N_{\rm eff}$ model and determine regions of the parameter space that can simultaneously resolve the $H_{0}$ and $S_{8}$ tensions inprep . The likelihood analysis will also include extra free parameters, e.g., the sum of neutrino masses: $\Lambda_{s}{\rm CDM}+N_{\rm eff}+\sum m_{\nu}$.
@@ Acknowledgments
We thank Niccolò Cribiori and Eleonora Di Valentino for valuable discussions. The research of LAA is supported by the U.S. National Science Foundation (NSF grant PHY-2112527). The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland.

### v4

@@ Abstract
Very recently, an empirical conjecture to simultaneously resolve the $H_{0}$ and $S_{8}$ tensions has been contemplated Akarsu:2023mfb . The conjecture, which is rooted on the graduated dark energy model Akarsu:2019hmw ; Akarsu:2021fol ; Akarsu:2022typ , postulates that $\Lambda$ may have switched sign (from negative to positive) at critical redshift $z_{c}\sim 2$. The so-called “$\Lambda_{s}$ model” then suggests that the Universe may have recently experienced a phase of rapid transition from anti-de Sitter (AdS) vacuum to de Sitter (dS) vacuum. However, the AdS swampland conjecture Lust:2019zwm indicates that AdS vacua are an infinite distance apart from dS vacua in moduli space. In fact, the AdS distance conjecture, taken at face value, implies that it is not possible to cross the barrier $\Lambda=0$ and to have a transition from a negative to a positive cosmological constant. This follows because the distance in the space of metrics in terms of $\Lambda=0$ is proportional to $-\log|\Lambda|$, and this expression diverges as $\Lambda\rightarrow 0$. However, we note that this no-go theorem is valid at zero temperature, where the number of light particles is (in general) constant. At finite temperature, particles can decay and hence the number of light particles can change. In this way the minima of the potential can be lifted. All in all, the postulated switch in sign of $\Lambda$ at zero temperature seems, at first sight, unlikely. In this Letter we investigate whether quantum effects derived from the Casimir energy of fields inhabiting the bulk could produce the hypothesized AdS $\to$ dS crossover transition. [^1] For related work, see Ye:2020btb .
Before proceeding, we pause to discuss the benefits of the $\Lambda_{s}$ model. The model simultaneously resolves the major cosmological tensions including: $H_{0}$, $S_{8}$, and SNe Ia absolute magnitude Akarsu:2023mfb . In addition, it can accommodate the BAO Lyman-$\alpha$ discrepancy Akarsu:2019hmw . [^2] The recent investigation presented in Akarsu:2023mfb is based on: the Planck CMB data Planck:2019nip , the Pantheon+ SNe Ia sample Brout:2022vxf , the data release of KiDS-1000 KiDS:2020suj , and the (angular) transversal 2D BAO data on the shell Nunes:2020hzy ; deCarvalho:2021azj , which are less model dependent than the 3D BAO data used in previous studies of $\Lambda_{s}$. It is important to stress that the BAO 3D data sample assumes $\Lambda$CDM to determine the distance to the spherical shell, and hence could potentially introduce a bias when analyzing beyond $\Lambda$CDM models Bernui:2023byc . Furthermore, for flat cosmology, $\Omega_{m}+\Omega_{\Lambda}\simeq 1$, from which it follows that the mean energy density in the universe is equal to the critical density, where $\Omega_{m}$ is the fractional nonrelativistic matter density and $\Omega_{\Lambda}$ is the cosmological-constant energy density. This implies that $\Omega_{m}>1$ when $\Lambda<0$, in agreement with the otherwise puzzling JWST observations Adil:2023ara . [^3] See however Sabti:2023xwo and Padmanabhan:2023esp .
@@
This potential possesses a minimum at $R=R_{1}$, for which one can show that it always satisfies $R_{1}>R_{0}$. So the radius gets increased due to the uplift, and hence the associated species scale Dvali:2007hz ; Dvali:2007wp ; vandeHeisteeg:2022btw ; Cribiori:2022nke ; vandeHeisteeg:2023dlw decreases (see also the discussion below). The species entropy was introduced in Cribiori:2023ffn , and it is defined in four space-time dimensions as $S_{\rm sp}=M_{*}^{-2}$. Following the law of thermodynamics, it is also argued in Cribiori:2023ffn that any dynamical motion in moduli space should always follow a path, for which $S_{\rm sp}$ does not decrease. With $M_{*}\simeq R^{-1/3}$, it follows that species thermodynamics allows only for transition processes where the radius of the extra dimension increases (or stays constant) due to the uplift.
@@
Note that for $a-a^{c}\to 0^{+}$, the minimum $R_{1}$ is near $R_{0}$, consistently with the quadratic approximation given in (1). Thus, the parametric dependence between the vacuum energy and the radius at its minimum is the same as before the uplift, in accordance with the dark dimension scenario Montero:2022prj . A point worth noting at this juncture is that the relation $|V|=R^{-4}$ is at the heart of vacuum energy calculations in string theory, see Anchordoqui:2023laz for a recent discussion.
@@
In the spirit of Schwarz:2024tet , we model the dark dimension as a space with two boundaries: an interval with end-of-the-world 9-branes attached at each end. After a suitable compactification of six dimensions the space boundaries can be approximated by 3-branes, and one of them can be identified with the Standard Model 3-brane which portrays our non-compact 4D world, at least up to energies ${\cal O}(10~{}{\rm TeV})$ and perhaps much further. The line interval along the dark dimension can also be understood as a semicircular dimension endowed with $S^{1}/\mathbb{Z}_{2}$ symmetry. [^6] In the context of toroidal orientifold models compact dimensions are associated to $S^{1}/\mathbb{Z}_{2}$ symmetry which is needed for obtaining a 4D chiral model ($S^{1}$ does not lead to chirality).
@@
where $\Lambda_{5}$ is the 5D cosmological constant, $r\equiv\langle R\rangle$, $T_{4}$ is the total brane tension, and
$V_{C}(R)=\pi R\left({r\over R}\right)^{2}\sum_{i}(-1)^{s_{i}}\ N_{i}\ \rho(R,m)$ (9)
corresponds to the Casimir energy, and where the sum goes over all 5D states in the spectrum, $N_{i}$ is the number of degrees of freedom of the $i$-th particle, $s_{i}=0\,(1)$ for bosons (fermions), and the factor $(r/R)^{2}$ comes from the rescaling of the 4D metric in the transformation to the Einstein frame Anchordoqui:2023etp . The Casimir energy density for a particle of mass $m$ is given by
@@
where $K_{5/2}(x)$ is the Bessel function Arkani-Hamed:2007ryu . [^7] The Casimir energy per unit of volume is the same for Dirichlet and Neumann boundary conditions Barone:2003rn .
@@
$V_{C}(R)=\sum_{i}\frac{\pi r^{2}}{32\pi^{7}R^{6}}\ (N_{F}-N_{B})\ \Theta(R_{i}-R)\,,$ (12)
where $m_{i}=R_{i}^{-1}$, $\Theta$ is a step function, $N_{F}-N_{B}$ stands for the difference between the number of light fermionic and bosonic degrees of freedom Anchordoqui:2023etp . Eventually, the fermionic degrees of freedom overwhelm the bosonic contribution, giving rise to possible minima, as long as $R_{i}<R_{\rm max}$.
Figure 1: The potential $V(R)$ for $\Lambda_{5}=(22.6~{}{\rm meV})^{5}$ and $|T_{4}|=(24.2~{}{\rm meV})^{4}$ considering $N_{F}-N_{B}=3$ (AdS) and $N_{F}-N_{B}=7$ (dS).
To get further insights into this idea we envision a next-to-minimal hidden sector to characterize the deep infrared region. It consists of: the 5D graviton (contributing with $N_{B}=5$), a massive gauge boson ($N_{B}=4$), and three massive bulk fermions (each contributing with $N_{F}=4$). These fermions could be identified with sterile neutrinos furnished by Dirac bulk masses, which weaken the bounds on the size of the fifth dimension from neutrino oscillation data Anchordoqui:2023wkm . Altogether, this gives a difference $N_{F}-N_{B}=3$. Now, we would further assume that the gauge boson is unstable and decays into two neutrinos. The 5D gauge coupling has dimension of mass-1/2, which we can write as $g/M_{*}^{1/2}$ where $M_{*}$ is the 5D gravity scale (of order the string scale) and $g$ a dimensionless parameter. The total decay width is then given by
@@
We can also envision a more complex dark sector in which the gauge field is replaced by a complex scalar with Yukawa coupling to the neutrinos (that could provide them with mass). The spectrum has addition of an extra complex scalar, such that before its decay $N_{F}-N_{B}=5$ and after decay $N_{F}-N_{B}=7$. In Fig. 2 we show the corresponding AdS $\to$ dS transition for this scenario. For this scenario, $\Delta N_{\rm eff}=0.27$.
Figure 2: The potential $V(R)$ for $\Lambda_{5}=(22.6~{}{\rm meV})^{5}$ and $|T_{4}|=(24.2~{}{\rm meV})^{4}$ considering $N_{F}-N_{B}=5$ (AdS) and $N_{F}-N_{B}=7$ (dS).
The next-to-minimal dark sectors that can accommodate the AdS $\to$ dS crossover transition can be nicely combined with the dark dimension proposal to elucidate the radiative stability of the cosmological hierarchy $\Lambda/M_{p}^{4}\sim 10^{-120}$ Montero:2022prj . As can be seen in Figs. 1 and 2, the connection is possible for a compactification radius $R_{\perp}^{\rm(AdS)}\sim 8~{}\mu{\rm m}$, which implies the Kaluza-Klein (KK) tower of the dark dimension opens up at $m_{\rm KK}^{\rm(AdS)}\sim 155~{}{\rm meV}$. After the vacuum undergoes the AdS $\to$ dS transition the compact space slightly enlarges to $R_{\perp}^{\rm(dS)}\sim 9~{}\mu{\rm m}$ and so $m_{\rm KK}^{\rm(dS)}\sim 138~{}{\rm meV}$. This in turn makes a shift of the species scale,
@@
Another possible course of action to accommodate an AdS $\to$ dS transition (with $N_{F}-N_{B}=6$ in the AdS phase and $N_{F}-N_{B}=7$ in the dS phase) is to assume that a real scalar $\varphi$ has a potential with two local minima with very small difference in vacuum energy and bigger curvature (mass) of the lower one. More concretely, the potential is given by
@@
where $V_{\varphi}=\partial V/\partial\varphi$. The fourth degree of freedom is adjusted by requiring the height of the barrier to accommodate a decay rate of the false vacuum to be
@@
where $Z(T)=\int[dx]\ e^{-S_{E}/\hslash}$ is the path integral, $[dx]$ is a functional measure, and $S_{E}$ the Euclidean action Coleman:1980aw ; Hawking:1981fz . Finally, we take $\Lambda_{5}\sim(22.6~{}{\rm meV})^{5}$ and $|T_{4}|=(24.2~{}{\rm meV})^{4}$ to obtain the vacuum energy densities of the AdS and dS phases shown in Fig. 3. For this scenario, $\Delta N_{\rm eff}=0.25$.
Figure 3: The potential $V(R)$ for $\Lambda_{5}=(22.6~{}{\rm meV})^{5}$ and $|T_{4}|=(24.2~{}{\rm meV})^{4}$ considering $N_{F}-N_{B}=6$ (AdS) and $N_{F}-N_{B}=7$ (dS).
At this stage, it is constructive to connect with contrasting and complementary perspectives to comment on two caveats of our analysis: (i) although the cosmological models are motivated by swampland constraints and string constructions, they have no concrete UV completion; (ii) another aspect of this analysis which may seem discrepant at first blush is the fact that the (one loop) Casimir energy is a zero-$T$ effect. However, in the example of the real scalar (which not unexpectedly is at zero temperature) the argument to understand the transition is essentially the same than the one in finite temperature models because the number of light degrees of freedom changes due to a different transition of the 5D scalar field. In other words, the model avoids finite temperature requirements and relies on an ordinary vacuum decay in five dimensions. This obviously implies that the AdS vacuum is not a true vacuum. The vacuum in the radius modulus is determined by the contribution to the Casimir potential of the number of light degrees of freedom. This number changes discontinuously due to an ordinary vacuum decay of a 5D scalar field which satisfies the swampland AdS conjecture (for instance it can be dS to dS). This change drives the AdS to dS transition in the radius modulus, which is therefore discontinuous as in first order transitions.
It is of interest to further investigate the impact of adding extra-relativistic degrees of freedom into the $\Lambda_{s}{\rm CDM}$ model. In particular, a value of $\Delta N_{\rm eff}\neq 0$ would accelerate the universe before the CMB epoch, and hence could modify the critical redshift $z_{c}$ for the AdS $\to$ dS crossover transition. A study along this line is presented in an accompanying paper Anchordoqui:2024gfa where we use the Boltzmann solver CLASS Blas:2011rf ; Lesgourgues:2011rh in combination with MontePython Audren:2012wb ; Brinckmann:2018cvx (and the latest results from the Planck mission Planck:2019nip , Pantheon+ SNe Ia Brout:2022vxf , BAO Nunes:2020hzy ; deCarvalho:2021azj , and KiDS-1000 KiDS:2020suj ) to perform a Monte Carlo Markov Chain study of the $\Lambda_{s}{\rm CDM}+N_{\rm eff}$ model and determine regions of the parameter space that can simultaneously resolve the $H_{0}$ and $S_{8}$ tensions.
@@ Acknowledgments
We thank Niccolò Cribiori, Eleonora Di Valentino, and Arthur Hebecker for valuable discussions. The research of LAA is supported by the U.S. National Science Foundation (NSF grant PHY-2112527). The work of D.L. is supported by the Origins Excellence Cluster and by the German-Israel-Project (DIP) on Holography and the Swampland.